Diamond. Thanks for the help. I will comment Diamond. Thanks for the help. I will comment if the answer is INCORRECT in hopes that you can retry the problem before I give a negative rating.
Answer #1

The image of diamond is virtual.

Now from lens equation for focal length \frac{1}{f}=\frac{1}{v}-\frac{1}{u}

-\frac{1}{5.05}=\frac{1}{v}+\frac{1}{10.9}

  -\frac{1}{5.05}-\frac{1}{10.9}=\frac{1}{v}

Therefore image position is v= -3.45cm

from definition of magnification m = \frac{h_{i}}{h_{o}}= \frac{v}{u}

The height of image is hi = 3.45x3.21/10.9 = 1.02cm